Cavalier Copper shows on its most recent balance sheet $840 million in total liabilities, $10 billion...

  1. Cavalier Copper shows on its most recent balance sheet $840 million in total liabilities, $10 billion in treasury stock (recorded at par of $1/share) and $520 million in common stock outstanding (also recorded at par of $1/share). The company had negative retained earnings of $40 million at the beginning of its fiscal year, and it made a net profit of $60 million during the year. The company’s stock sells for $4/share on the date of the balance sheet. The company’s total assets are.
    1. $1.38 billion
    2. $10.84 billion
    3. $1.46 billion
    4. $2.94 billion
  2. Using the same information as in question #28 and assuming the company discloses an additional $100 million in operating lease commitments with a net present value of $100 million, which it treats as debt, the company’s total debt to total capital ratio should be.
    1. .54
    2. 1.74
    3. .45
    4. 1.56
Answer #1

Ans 1)total assets= total liabilities

total Liabilities = liabilities +shareholders fund

shareholders fund=common stock +retained earning as on the balance sheet date

retained earnings as on bs date= profit at the year-end- loss at the beginning of the year

=60-40= 20 million

shareholders fund =520+20

=540

total liabilities=840+540=1380million i.e 1.38 billion

therefore total asset=1.38 billion

Therefore option A is correct

Ans 2)100 million addition to total liabilities =840+100=940 million

debt/equity ratio=debt/equity

=940/540=1.74

therefore option B is correct